OpenStudy (daniella1):

math computation a pack of biscuits has a mass of 1/3 kg,jason eat 2/3 of the biscuits,. what is the mass of biscuits left.

OpenStudy (kevin):

\[\frac{ 1 }{ 3 } - (\frac{ 2 }{ 3 } \times \frac{ 1 }{ 3 })\] \[= \frac{ 1 }{ 3 } - \frac{ 2 }{ 9 }\] \[=\frac{ 1 }{ 9 }\]